My question is: Cube root of (54/9) minus 3 times the square root of 1296.
Answer by checkley71(8403)

CUBERT54/9=3CUBERT2/9-3SQRT1296=3CUBERT2/9-3*36=3(CUBERT2/9-36)M ANSWER.